What is the effective tax rate for an investor who owns a property which has a before-tax yield of 12.30% and an after-tax yield of 10.12%? (Rounded to nearest whole %)

1 35%

2 18%

3 25%

4. 15%
Answer #1

Before tax yield*(1-tax rate)=Aftertax yield

12.3*(1-tax rate)=10.12

1-tax rate=(10.12/12.3)

tax rate=1-(10.12/12.3)

=18%(Approx).